OpenStudy (anonymous):

An object is launched upward at 64 ft/sec from a platform that is 80 feet high. What is the objects maximum height if the equation of height (h) in terms of time (t) of the object is given by h(t) = -16t2 + 64t + 80?

OpenStudy (anonymous):

h(t) = -16t^2 + 64t + 80 = 0 The object's trajectory would be a parabola, with origine O as the starting point. The max. of the parabola occurs at t = -b/2a (axis of symmetry of the parabola): t = -64/-32 = 2. At this moment t = 2, the height of the object is: h(2) = -16 (2)^2 + 64*(2) + 80 = .....

OpenStudy (anonymous):

thank you so much